OpenStudy (anonymous):

The formula for voltage (in volts) is given by V = IR, where I is the current in amperes, and R is the resistance in ohms. If a light bulb uses 0.4 amperes of current at a voltage of 1.8 volts, the formula to find R is_____ , and the resistance of the bulb is_____ ohms.

OpenStudy (anonymous):

Answers for 1st: R= I/V R=V/I R=VI

OpenStudy (anonymous):

Answers for 2nd: 4 4.5 5

OpenStudy (anonymous):

Do you know how to get R by itself in the equation?

OpenStudy (anonymous):

would you divide out the I?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so R= V/I?

OpenStudy (anonymous):

yes. and then just plug in the numbers for the 2nd part

OpenStudy (anonymous):

4.5?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

THANK YOU!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

medal, fan and testimonial heading your way

OpenStudy (anonymous):

thank you